OpenStudy (ttop0816):

Please help!! Find the center and radius of the given circle and graph it. x^2+y^2+10y+21=0 I got how to set up the equation but where is a pair of x value for x^2 when there is 10y for y^2? :(

OpenStudy (ttop0816):

@phi please help if you are not busy! :)

OpenStudy (phi):

do you know how to "complete the square" on the y part ?

OpenStudy (ttop0816):

x^2+y^2+10y=-21 (x^2...??)+(y+2+10y....)=-21 for y, 10/2=5^2=25 (x^2...??)+(y+2+10y+25)=-21+25

OpenStudy (ttop0816):

I got it till here

OpenStudy (phi):

I assume you mean (y^2 +10y+25) (not y+2 + 10y +25)

OpenStudy (phi):

anyway, the point of completing the square is so we can write (y^2 +10y+25) as (y+5)^2 thus your equation is now \[ x^2 + (y+5)^2 = 4 \]

OpenStudy (phi):

next, match that up with the "generic" equation for a circle (x-a)^2 + (y-b)^2 = r^2 the center of the circle will be at (a,b) and the radius is r

OpenStudy (ttop0816):

so the circle would be (0,5) radius would be 16??

OpenStudy (phi):

your equation looks a bit like the generic equation , but we have to tweak it to make it look the same form x^2 + (y+5)^2 = 4 (x-a)^2 + (y-b)^2 = r^2 looking at the right-hand side, we want 4 to match r^2 we know 4 is 2*2 or 2^2. so we would write it this way x^2 + (y+5)^2 = 2^2 and now 2^2 matches r^2 and we see what r is: it is 2

OpenStudy (ttop0816):

ohhh I see, then would the circle be correct?

OpenStudy (phi):

next, we don't have y- b matching with y+5 but we could write y+5 as y - (-5) that means the same thing, but now we have y - a number and we can match x^2 + (y- -5)^2 = 2^2 (x-a)^2 + (y-b)^2 = r^2 and we see b matches up with -5

OpenStudy (phi):

finally , we can write x as (x-0) which is still "x", but it now can match up with the generic equation (x-0)^2 + (y- -5)^2 = 2^2 and we see the center is (0,-5) and the radius is 2

OpenStudy (phi):

is it not clear ?

OpenStudy (ttop0816):

Sorry I was just trying out what you said on my paper. Yes!! I totally get it now

OpenStudy (ttop0816):

If you don't mind, could you help with this question as well? Find the slope of the line that is a parallel and perpendicular to the given line. 3x+y=(3x+2)

OpenStudy (ttop0816):

I know how to find both ways but I'm just confused because the slope is zero. y=3x-3x+2 y=2?

OpenStudy (phi):

to do this problem, notice the trick we use we change y+5 to y- -5 so that we can match that to y-b the other trick was to write x as x-0 though if you do enough problems you will remember x^2 by itself means the center of the circle will have 0 for its x value. and you will remember that y+5 means the y value of the center is -5 if you don't do a lot of these problems, then you have to remember the tricks.

OpenStudy (phi):

3x+y=(3x+2) yes that simplifies to y=2 y=2 means points like (0,2), (1,2), (2,2) all lie on that line if you plot those points you see a line going horizontal (or as you noted, its slope is 0, and that means a line going sideways) they want a line parallel to y=2 pick slope=0 but use a different y intercept.

OpenStudy (phi):

they also want a perpendicular line to y=2 that means you want a line that forms a right angle i.e. 90 degrees that would be a straight up/down vertical line do you know the equation of a vertical line ? you write x= 0 or x=1 or x=2 all of those are vertical lines

OpenStudy (phi):

the slope of a vertical line is called "undefined" reading the question, they want the slope of a parallel and perpendicular line the answer is : 0 and "undefined"

OpenStudy (ttop0816):

Thanks!